# STCast: Adaptive Boundary Alignment for Global and Regional Weather Forecasting

### v1

@@ Introduction
Extreme Events Assessment. Extreme weather events, particularly tropical cyclones, pose significant societal risks, necessitating accurate prediction capabilities (Wang et al. 2025a, b). To evaluate our method’s performance under such critical conditions, we analyze two recent typhoon events: Typhoon Ewiniar (May 2024) and Typhoon Yinxing (November 2024) (Ying et al. 2014; Lu et al. 2021). As visualized in Fig. 9d-e, STCast’s 72-hour track forecasts demonstrate substantially closer alignment with observed paths compared to operational ECMWF predictions for both systems. This improved track accuracy highlights STCast’s enhanced capability for extreme event forecasting. Complementary visualization in Fig. 9a-c further dissects Typhoon Yinxing’s evolution, contrasting initial conditions, STCast predictions, and ground-truth. Our model consistently captures the cyclone’s structural development and translational dynamics, validating its physical representation of intense meteorological systems. More numerical comparison is shown in Appendix.
More Competitors. ClimaX (Nguyen et al. 2023), EWMoE (Gan et al. 2025), Keisler (Keisler 2022), Stormer (Nguyen et al. 2025), VAMoE (Chen et al. 2025), FourCastNet (Kurth et al. 2023), ClimODE (Verma, Heinonen, and Garg 2024), WeatherGFT (Xu et al. 2024), and GenCast (Price et al. 2025) are shown in Appendix.
Computation comparisons, additional visualization, and additional results are provided in Appendix.
Ablation Study
@@
To further verify the effectiveness of each proposed module and strategy, we conduct comprehensive ablation studies reported in Tab. 2. The experiments are split into two groups. High-resolution regional forecasting: (1) STCast w/o SAA: we remove the SAA module and follow the same protocol as Oneforecast to predict regional variables; (2) STCast w/o Global-Regional Distribution: we discard the global-regional distribution initialization in SAA and instead use random initialization for the global-regional correlation; (3) STCast w SAA: the complete STCast model. Low-resolution global forecasting: (4) STCast w/o TMoE: we replace the Temporal Mixture-of-Experts (TMoE) with MLP block; (5) STCast w/o Month Embedding: we remove the month embedding from TMoE and fall back to a classical Mixture-of-Experts; (6) STCast w TMoE: the complete STCast model. Comparisons among (1)–(3) and (4)–(6) reveal that removing any component consistently degrades performance on both regional and global tasks. While the absence of SAA or TMoE causes noticeable drops, the most substantial drops occur when eliminating global-regional distribution (regional: +0.22 RMSE in 10-day) and month embedding (global: +0.13 RMSE in 10-day). These results confirm the critical roles of every component and setting in enhancing the overall effectiveness of STCast.
Conclusion
@@
In this work, we introduce an adaptive attention map within the Spatial-Aligned Attention (SAA) module to provide dynamic boundary conditions for regional forecasting. Beyond regional task, we embed a Temporal Mixture-of-Experts (TMoE) into the Spatial-Temporal Forecasting (STCast), casting weather prediction as a multi-task problem and delegating monthly sub-tasks to specialized experts. Consequently, STCast simultaneously addresses 4 distinct challenges: low-resolution global forecasting, high-resolution regional forecasting, extreme-event assessment, and ensemble weather forecasting. Comprehensive experiments and ablation studies confirm that STCast consistently outperforms competing methods across all evaluated scenarios.
Appendix
STCast: Adaptive Boundary Alignment for Global and Regional
Weather Forecasting
Dataset Details
Dataset and Baselines
In this work, we conduct experiments on a popular weather dataset, i.e., ERA5111https://cds.climate.copernicus.eu/ (Hersbach et al. 2020), provided by the ECMWF (Molteni et al. 1996). As shown in Tab. 4, ERA5 dataset is a reanalysis atmospheric dataset, consisting of the atmospheric variables from 1979 to the present day with a 0.25°spatial resolution with . The atmospheric variables include 5 upper-air variables (Z, Q, U, V, T) on 13 levels and 5 surface variables (T2M, U10, U10, MSL, SP). The low-resolution global forecasting is trained on 70 variables () with 40 years atmospheric dataset from 1979 to 2019 with 1.4°spatial resolution. At the same time, the high-resolution regional weather forecasting task is trained in the same period with 5 surface variables on Eastern Asia (7.5°W114°E-36°W172.5°E) with 0.25°spatial resolution.
Data processing
To address disparities among variables, all model inputs are normalized to ensure consistency. Using the training dataset spanning 1979–2019, we compute the mean and standard deviation for each variable. Normalization is then performed by subtracting the respective mean and dividing by the corresponding standard deviation.
Implementation Details
The main structure of this work follows the backbone of Flash Attention (Dao et al. 2022). We apply the AdamW optimizer with 0.0002 learning rate to the model training. In both global and regional forecasting tasks, we train 100 epochs and set batch size to 16. Our model is trained with PyTorch using 16 NVIDIA Tesla A100 GPUs. In the global forecasting stage, we train the whole model. While in the regional forecasting stage, we only need to train the Spatial-Aligned Attention(SAA) module and freeze the main structure. More implementation details are provided in the logs.
Code Available
We provide some code in the Supplementary Materials for our STCast, OneForecast, GraphCast and related baselines. Baselines are collected from its respective official GitHub repository.
Additional Related Works
Time-series Methods
Before the emergence of recent data-driven forecasting methods on large-scale atmospheric datasets such as ERA5 (Hersbach et al. 2020), several time-series approaches had already been applied to weather forecasting. These earlier works typically framed regional forecasting as a video prediction task, employing spatio-temporal convolutional layers or Transformer blocks to process the input data. For instance, SimVP (Gao et al. 2022) and PastNet (Wu et al. 2024b) employed spatio-temporal convolutions to forecast regional atmospheric images, while STTN (Ji et al. 2024) and PKD-STTN (He, Ji, and Lei 2024) utilized spatio-temporal Transformer blocks. Concurrently, HiSTGNN (Ma et al. 2023) and STCWF (Gong et al. 2024) introduced spatio-temporal graph neural networks and contrastive learning, respectively, to the weather forecasting domain. Notably, PastNet (Wu et al. 2024b) further distinguished itself by incorporating physical principles into its neural network architecture.
Although the aforementioned time-series forecasting methods are commonly referred to as spatio-temporal approaches utilizing spatio-temporal neural networks (NNs), our proposed STCast fundamentally differs from them in several key aspects: (1) Motivation: Previous works aim to capture implicit temporal correlations among time-series inputs and spatial dependencies within the input domain using neural components such as convolutional layers. In contrast, STCast explicitly models temporal correlations across monthly atmospheric patterns and geographical relationships across the global domain. (2) Architecture: Prior works typically model spatial and temporal dimensions through convolution layers or Transformer blocks, where spatial modeling is performed via convolution kernels applied to image grids, and temporal modeling is achieved by extending these kernels across time steps. In contrast, STCast introduces a novel spatio-temporal modeling framework based on monthly Gaussian distributions and global–regional representations, enabling more structured and interpretable learning across both spatial and temporal domains. (3) Benchmarks: Unlike previous time-series methods that are typically evaluated on small-scale, low-resolution datasets such as SEVIR (Veillette, Samsi, and Mattioli 2020) and WD (Wang et al. 2019), STCast is the first to explore explicit spatio-temporal correlations within a realistic Earth system. It is evaluated on the ERA5 dataset (Hersbach et al. 2020), demonstrating its scalability and effectiveness in large-scale global weather forecasting.
| Name | Description | Levels | Resolution | Lat-Lon Range | Time |
| Low-resolution Global Weather Forecasting | |||||
| Z | Geopotential | 13 | -90°S180°W-90°N180°E | 1979-2020 | |
| Q | Specific humidity | 13 | -90°S180°W-90°N180°E | 1979-2020 | |
| U | x-direction wind | 13 | -90°S180°W-90°N180°E | 1979-2020 | |
| V | y-direction wind | 13 | -90°S180°W-90°N180°E | 1979-2020 | |
| T | Temperature | 13 | -90°S180°W-90°N180°E | 1979-2020 | |
| t2m | Temperature at 2m height | Single | -90°S180°W-90°N180°E | 1979-2020 | |
| u10 | x-direction wind at 10m height | Single | -90°S180°W-90°N180°E | 1979-2020 | |
| v10 | y-direction wind at 10m height | Single | -90°S180°W-90°N180°E | 1979-2020 | |
| msl | Mean sea-level pressure | Single | -90°S180°W-90°N180°E | 1979-2020 | |
| sp | Surface pressure | Single | -90°S180°W-90°N180°E | 1979-2020 | |
| High-resolution Regional Weather Forecasting | |||||
| t2m | Temperature at 2m height | Single | 7.5°S114°W-36°N172.5°E | 1979-2020 | |
| u10 | x-direction wind at 10m height | Single | 7.5°S114°W-36°N172.5°E | 1979-2020 | |
| v10 | y-direction wind at 10m height | Single | 7.5°S114°W-36°N172.5°E | 1979-2020 | |
| msl | Mean sea-level pressure | Single | 7.5°S114°W-36°N172.5°E | 1979-2020 | |
| sp | Surface pressure | Single | 7.5°S114°W-36°N172.5°E | 1979-2020 | |
Model Details
The global weather forecasting framework comprises three core components: Encoder, Processor, and Decoder, and two supplementary settings: Reconstruction and Loss Function. In this work, the Encoder and Decoder implementations follow Flash-Attention (Dao et al. 2022), while the Reconstruction setting and Loss Function design adopt the same setting of VA-MoE (Chen et al. 2025). For the forecasting task, the AI model predicts future atmospheric states from historical fields as . Detailed configurations for all five elements are provided below.
Encoder
Atmospheric variables across pressure levels are organized into a 3D tensor , where and denote the global grid height and width, respectively, and is the number of variables. This tensor is projected into patch embeddings via a convolutional layer Conv with stride equal to the patch size:
where we set .
In addition to patch embedding, we incorporate a learnable positional embedding matrix, denoted as , to encode spatial information within the Encoder module. This matrix is initialized using a truncated normal distribution defined as:
where and represent the mean and standard deviation, and and denote the lower and upper bounds of the distribution, respectively. The positional embedding is subsequently added to the input features prior to processing, yielding the updated representation .
Processor
The processor consists of a sequence of Transformer blocks, each comprising multi-head attention, a TMoE (Temporal Mixture-of-Experts) module, layer normalization, and residual connections. The operations within a single block can be formally expressed as:
where Attention denotes the attention mechanism, LN represents layer normalization, and TMoE refers to the Temporal Mixture-of-Experts module as described in the main paper.
Following the design principles of FlashAttention (Dao et al. 2022) and VA-MoE (Chen et al. 2025), we adopt an alternating strategy that combines window-based attention and global self-attention. This hybrid approach enables the model to effectively capture both local and global dependencies in the input distribution.
Decoder
The Decoder module in this work is implemented as a multi-layer perceptron (MLP), which predicts the atmospheric variables for the next timestep. The decoding operation is defined as:
The MLP consists of two linear layers separated by a non-linear activation function. Specifically, the decoding process can be expressed as:
where GELU denotes the Gaussian Error Linear Unit activation function. This structure enables the decoder to model complex relationships in the input features and generate accurate predictions for the subsequent timestep.
Reconstruction
To ensure training stability, we introduce an auxiliary reconstruction path that directly connects the Encoder and Decoder modules to reconstruct the input variables. This design complements the primary prediction path, which consists of the Encoder, Processor, and Decoder modules. Notably, the Encoder and Decoder are shared across both paths. The overall process is defined as:
where Encoder, Processor, and Decoder denote the respective modules in our framework. The reconstruction path facilitates the learning of robust representations by encouraging the model to preserve essential input information throughout the encoding and decoding stages.
Under this configuration, the Encoder and Decoder modules are dedicated to encoding and decoding the input variables, respectively, while the Processor is solely responsible for prediction. By excluding the Processor from the encoding and decoding stages, the framework avoids unnecessary computational overhead, thereby enhancing efficiency without compromising performance.
Loss function
To address both prediction and reconstruction tasks, we employ the L2 loss function to quantify point-wise errors between the predicted outputs and the ground truth. The prediction loss and reconstruction loss are defined as follows:
where and denote the reconstructed and predicted outputs, respectively. The operator Mean computes the average error across multiple dimensions. A weighting hyperparameter is introduced to balance the two objectives in the final loss function.
| Model | Params(M) | MACs(G) | GPUs | Training Time |
| Fengwu (Chen et al. 2023a) | 153.49 | 132.83 | 32 A100 | 17 days |
| FourCastNet (Kurth et al. 2023) | - | - | 64 A100 | 16 hrs |
| GraphCast (Lam et al. 2023) | 28.95 | 1639.26 | 32 TPUv4 | 4 weeks |
| Pangu-Weather (Bi et al. 2023) | 23.83 | 142.39 | 192 V100 | 64 days |
| VA-MoE (Chen et al. 2025) | 665.37 | - | 32 A100 | 6 days |
| OneForecast (Gao et al. 2025) | 24.76 | 509.27 | 16 A100∗ | 8 days∗ |
| Ours (STCast) | 616.16 | 436.12 | 16 A100 | 5 days |
Experiments Details
Evaluation Metric
In this work, we evaluate the forecasting performance between our STCast and other methods on RMSE (Root Mean Square Error) and ACC (Anomalous Correlation Coefficient), which can be defined as:
where and denotes the predicted variables and ground-truth at the horizontal coordinate and time ; and denote the length of latitude and longitude in the global region.
Considering the difference in the distribution of atmospheric variables at latitudes, we introduce the latitude-dependent function to weight the atmospheric variables. The function is formulated as:
where and denote the latitude at index and , respectively.
For typhoon track prediction, we evaluate model performance using two metrics: Mean Distance Error (MDE) and Haversine Distance. First, the Haversine Distance is computed between the predicted typhoon center and the ground-truth location to account for the curvature of the Earth. Subsequently, the MDE is used to quantify the average positional error across all time steps. These metrics are defined as follows:
| MDE | (19) | |||
where is the average radius of the earth, and are the latitude and longitude of two points in earth system. and are the latitude-longitude coordinates of predicted and real points.
Typhoon Tracking
Following previous AI-based approaches (Bi et al. 2023; Magnusson et al. 2021), we identify the eye of a tropical cyclone as the location of the local minimum in mean sea level pressure (MSLP). For tracking purposes, the temporal resolution is set to 6-hour intervals.
This study focuses on two extreme cyclones: Typhoon Ewiniar and Typhoon Yinxing. Typhoon Ewiniar formed east of Mindanao on May 24, 2024, traversed the Philippine Sea, and recurved northeastward over the Okinawa–Ryukyu region. Typhoon Yinxing developed east of Yap Island on November 4, 2024, crossed the Philippine Sea, and entered the South China Sea. The initial conditions for these cyclones are set at 00:00 UTC on May 24, 2024, and 00:00 UTC on November 4, 2024, respectively. Ground-truth and ECMWF are obtained from TCData222tcdata.typhoon.org.cn.
Additional Results
Efficiency Analysis
As shown in Tab. 5, we compare the number of parameters, multiply–accumulate operations (MACs), GPU usage, and training duration across six baseline models. Although STCast contains more parameters and MACs than GNN-based methods such as GraphCast and OneForecast, its overall computational cost remains significantly lower than that of previous models, particularly Fengwu, Pangu-Weather, and GraphCast. These results demonstrate that STCast achieves superior performance while maintaining comparable computational efficiency.
Additional Global Forecasting Analysis
As illustrated in Fig. 10, we compare STCast with several state-of-the-art models across forecasting horizons ranging from 6 to 120 hours, evaluated on four key atmospheric variables. Experimental results show that STCast performs comparably to VA-MoE and Stormer in predicting Z500 and U10, while outperforming all baselines in T850 and T2M. These findings highlight the effectiveness of STCast and its integrated TMoE architecture in global weather forecasting, demonstrating its ability to capture both temporal correlations and seasonal variability across diverse meteorological variables.
Additional Typhoon Track Prediction
In addition to the typhoon track visualizations presented in the main paper, we compare the Mean Distance Error (MDE, in kilometers) between ECMWF and STCast, as shown in Fig. 11. Experimental results indicate that STCast achieves comparable performance to ECMWF in short-term forecasts, while significantly outperforming it in long-term predictions. Specifically, STCast yields track prediction errors of 67.10 km for Typhoon Yinxing and 109.34 km for Typhoon Ewiniar, substantially lower than ECMWF’s errors of 165.25 km and 138.82 km, respectively. These findings demonstrate the strong capability of STCast in extreme event assessment, particularly in accurately forecasting tropical cyclone trajectories over extended time horizons.
Additional Visualization
We provide more visualization about regional weather forecasting of 6-hour, 0.5-day, 1-day, 1.5-day, 2-day, 2.5-day, 3-day, 3.5-day, 4-day, 4.5-day, 5-day, 5.5-day, 6-day, 6.5-day, 7-day, 7.5-day, 8-day, 8.5-day, 9-day, 9.5-day, and 10-day in Fig. 12, Fig. 13, Fig. 14, Fig. 15, Fig. 16, Fig. 17, Fig. 18, Fig. 19, Fig. 20, Fig. 21, Fig. 22, Fig. 23, Fig. 24, Fig. 25, Fig. 26, Fig. 27, Fig. 28, Fig. 29, Fig. 30, Fig. 31, and Fig. 32.
We also provide more visualization about global weather forecasting of 6-hour, 0.5-day, 1-day, 1.5-day, 2-day, 2.5-day, 3-day, 3.5-day, 4-day, 4.5-day, 5-day, 5.5-day, 6-day, 6.5-day, 7-day, 7.5-day, 8-day, 8.5-day, 9-day, 9.5-day, and 10-day in Fig. 33, Fig. 34, Fig. 35, Fig. 36, Fig. 37, Fig. 38, Fig. 39, Fig. 40, Fig. 41, Fig. 42, Fig. 43, Fig. 44, Fig. 45, Fig. 46, Fig. 47, Fig. 48, Fig. 49, Fig. 50, Fig. 51, Fig. 52, and Fig. 53.

### v2

@@ 4.2 Main Results
4.3 Ablation Study
@@
To further verify the effectiveness of the proposed modules, we conduct comprehensive ablation studies reported in Tab. 2. The experiments are split into two groups. High-resolution regional forecasts: (1) STCast w/o SAA: we remove the SAA module and follow the same protocol as OneForecast to predict regional variables; (2) STCast w/o Global-Regional Distribution: we discard the global-regional distribution initialization in SAA and instead use Xavier uniform initialization for the global-regional correlation; (3) STCast w SAA: the complete STCast. Low-resolution global forecasts: (4) STCast w/o TMoE: we replace the Temporal Mixture-of-Experts (TMoE) with MLP block; (5) STCast w/o Month Embedding: we remove the month embedding from TMoE and fall back to a classical Mixture-of-Experts; (6) STCast w TMoE: the complete STCast. Comparisons among (1)–(3) and (4)–(6) reveal that removing any component consistently degrades performance on both regional and global tasks. While the absence of SAA or TMoE causes noticeable drops, the most substantial drops occur when eliminating global-regional distribution (regional: +0.22 RMSE in 10-day) and month embedding (global: +0.13 RMSE in 10-day). These results confirm the critical roles of every component and setting in enhancing the overall effectiveness of STCast.
TMoE. As illustrated in Tab. 4, we conduct ablation studies to evaluate the impact of experts’ numbers in TMoE: (1) 4 w G: 4 experts with quarterly allocation and Gaussian embedding; (2) 6 w G: 6 experts with two-month allocation and Gaussian embedding; (3) 12 w G: 12 experts with monthly allocation and Gaussian embedding; (4) 12 w/o G: 12 experts with monthly allocation and one-hot embedding. From settings (1) to (3), results show that increasing the number of experts improves performance, albeit with a significant rise in the parameters. Further increasing the experts’ number to 24, corresponding to a biweekly allocation scheme, results in the parameter of 1164.7M, which becomes computationally prohibitive. While finer temporal divisions yield better results, a balance between accuracy and efficiency is achieved by setting the experts’ number to 12. Comparing (3) and (4), we observe that Gaussian embeddings outperform one-hot embeddings in all settings. Unlike one-hot, which activate only one expert at a time, Gaussian enable the model to capture relationships both within and across months, resulting in better performance.
SAA. To evaluate the impact of distance metrics on the SAA, we performed an ablation study comparing three measures in Tab. 5: Manhattan, Euclidean, and Great Circle distances. The results show that while the choice of distance metric has a limited effect on overall performance, the Great Circle distance yields the lowest RMSE due to its superior accuracy in measuring distances on the Earth’s surface. The learned spatial distribution is illustrated in Fig. 3.
| Low-resolution Global Forecasts (RMSE) | ||||||
| Experts # | 6-hour | 1-day | 4-day | 7-day | 10-day | Params |
| 4 w G | 0.0750 | 0.1394 | 0.3147 | 0.5066 | 0.6433 | 484.9M |
| 6 w G | 0.0738 | 0.1329 | 0.2808 | 0.4555 | 0.5932 | 541.5M |
| 12 w G | 0.0617 | 0.1197 | 0.2578 | 0.4348 | 0.5763 | 654.8M |
| 12 w/o G | 0.0734 | 0.1351 | 0.2795 | 0.4509 | 0.5911 | 654.8M |
| High-resolution Regional Forecasts (RMSE) | |||||
| Distance Metric | 6-hour | 1-day | 4-day | 7-day | 10-day |
| Manhattan | 0.0493 | 0.0854 | 0.2068 | 0.3712 | 0.4945 |
| Euclidean | 0.0501 | 0.0862 | 0.2093 | 0.3892 | 0.5088 |
| Great Circle | 0.0478 | 0.0845 | 0.2051 | 0.3699 | 0.4921 |
5 Conclusion
@@
In this work, we introduce an adaptive attention map within the Spatial-Aligned Attention (SAA) module to provide dynamic boundary conditions for regional forecasting. Beyond regional task, we embed a Temporal Mixture-of-Experts (TMoE) into the Spatial-Temporal Forecasting (STCast), casting weather prediction as a multi-task problem and delegating monthly sub-tasks to specialized experts. Consequently, STCast simultaneously addresses 4 distinct challenges: low-resolution global forecasting, high-resolution regional forecasting, extreme-event assessment, and ensemble weather forecasting. Experiments and ablation studies confirm that STCast consistently outperforms competing methods across all evaluated scenarios.